# Stabilizing Circular Motion Within Nonconcentric Circular Boundary: A Möbius Transformation-Based Approach

### v1

@@ V-A Transformed Plane in Terms of Actual Plane Parameters
where we used the fact that ${\alpha(1-\alpha^{2})}/{|\alpha(1-\alpha^{2})|}=(\beta-\alpha)/|\beta-\alpha|=\Delta/|\Delta|$ after replacing 1 by the product $\alpha\beta$ and simplifying. Next, using (17) and (53), the inner product $\langle\rho,e^{i\gamma}\rangle$ can be readily expressed in terms of the actual plane parameters as $\mathcal{P}(r,\theta)$ in (49), and the error $\mathcal{E}$ in (19) can be written in terms of actual plane parameters as (50). Now, the result immediately follows from (26) and (47). ∎
@@ V-B Actual Plane in Terms of Transformed Plane Parameters
Please note that (54) has a singularity at $\rho=1$. However, we will prove in the following lemma that it is always well-defined in the framework of the considered problem.
@@
The validity of control (26) relies on the condition that $|\mathcal{E}(0)|<\delta_{T}$ must be satisfied in the transformed plane. It is necessary to map this requirement on the initial position and heading angle of the robot in the actual plane. In this direction, it follows from (50) that the condition
@@ VII-A Numerical Simulations
For illustration, consider the scenario described in Example 1, where we have nonconcentric circles with $\mu=\sqrt{5/2}$ and $\lambda=1/2$. Let the robot in the actual plane be characterized by its initial position and heading angle $[x(0),y(0)]^{\top}=[-0.9,-0.6653]^{\top}$ and $\theta(0)=-60^{\circ}$, respectively, and move with constant linear speed $v=1$ unit. It can be obtained that $|r(0)|=1.1192$ and $\phi(0)=-143.5274^{\circ}$. For $\alpha=\alpha_{s}=0.5$, $\delta_{T}=0.1325$ using (21). It can be verified that (66) is satisfied, since $\eta_{a}(0)=-0.045$, $\eta_{b}(0)=-0.0511$, and $\eta(0)=0.6428$, with $\operatorname{sgn}(\Delta)=1$ and $\sigma=0.5$, resulting in $(\eta_{a}^{2}(0)+\eta_{b}^{2}(0)=0.0046)<(\delta_{T}^{2}\eta^{2}(0)=0.0072)$. Using (17) and (39), the initial position and heading angle in the transformed plane are obtained as $[\rho_{x}(0),\rho_{y}(0)]^{\top}=[0.0016,-0.6039]^{\top}$ and $\gamma(0)=2.3326^{\circ}$, respectively. It can be further verified that $(|\mathcal{E}(0)|=0.106)<(\delta_{T}=0.1325)$.
@@
We conducted simulations using control law (6) with $\kappa=0.02$ and plotted various attributes in the actual plane as shown in Fig. 6. It is evident from Fig. 6(a) that the robot’s trajectory converges to the desired circular orbit $\mathcal{C}$ while being confined within the nonconcentric outer boundary $\mathcal{C}^{\prime}$. The absolute value of error (16) converges to the origin, as depicted in Fig. 6(b). The control $\omega$ with roots $\alpha_{s}$ and $\alpha_{\ell}$ is plotted in Fig. 6(c) and Fig. 6(d), respectively. For both values of $\alpha$, $\omega$ converges to $1$, as desired. Further, Fig. 7 illustrates the robot’s motion in the transformed plane and the results are as expected. Please note the difference in the trajectory plots Fig. 7(a) (trajectory-constraining) and Fig. 7(d) (obstacle-avoidance) obtained for $\alpha_{s}$ and $\alpha_{\ell}$, respectively. The robot’s motion is in the anticlockwise direction in Fig. 7(a), while it is in the clockwise direction in Fig. 7(d), in accordance with Theorem 3.
@@ VII-B Experimental Results
To assess the performance of the proposed controller (6) in real-world scenarios, we conducted experiments using a Khepera IV111k-team.com/khepera-iv differential drive ground robot. The experiment was facilitated by a motion capture (MoCap) validation system, providing precise feedback on the robot’s pose data, including position and heading angle. The MoCap system comprises Qualisys track manager (QTM) software for data recording and processing. Control commands were sent to the Khepera IV robot based on measurements obtained from QTM using the robot operating system (ROS).
To accommodate the size of Khepera IV robot, we adjusted the initial conditions slightly. The robot commenced its motion from the position $[x(0),y(0)]^{\top}=[-0.27\text{\leavevmode\nobreak\ m},-1.15\text{\leavevmode\nobreak\ m}]^{\top}$ with an initial heading angle of $\theta(0)=-30^{\circ}$. The robot was operated at a constant linear speed of $v=0.5$ m/s, adhering to its maximum speed of $0.814$ m/s as specified in the manual. Fig. 8(a) shows the top view of the experimental result at $t=21$ seconds where the red dotted line depicts the robot’s path. Throughout the experiment, the robot’s trajectory was recorded using the QTM software, as depicted in Fig.8(b). Here, the $X$-axis and $Y$-axis denote the real and imaginary axes, respectively. Further, Fig. 8(c) plots the control law applied during the experiment, demonstrating that the turn rate converges to $\omega\approx 0.5$ rad/sec at the steady state.
VIII Conclusion and Further Remarks
Appendix A
@@ 4. Case IV
Here, Case I and Case III refer to $\alpha=\alpha_{s}$ and the rest two cases are for $\alpha=\alpha_{\ell}$. This concludes the proof. ∎
Appendix B
@@
that, upon multiplication by ${\rm e}^{-i\phi}$ on both sides, yields
@@
where $\vartheta_{1}$ and $\vartheta_{2}$ can be simplified using (41) as

### v2

@@ V-A Transformed Plane in Terms of Actual Plane Parameters
where we used the fact that ${\alpha(1-\alpha^{2})}/{|\alpha(1-\alpha^{2})|}=(\beta-\alpha)/|\beta-\alpha|=\Delta/|\Delta|$ after replacing 1 by the product $\alpha\beta$ and simplifying. Next, using (17) and (53), the inner product $\langle\rho,{\rm e}^{i\gamma}\rangle$ can be readily expressed in terms of the actual plane parameters as $\mathcal{P}(r,\theta)$ in (49), and the error $\mathcal{E}$ in (19) can be written in terms of actual plane parameters as (50). Now, the result immediately follows from (26) and (47). ∎
@@ V-B Actual Plane in Terms of Transformed Plane Parameters
Note that (54) has a singularity at $\rho=1$. However, we will demonstrate in the following lemma that it remains well-defined within the context of the problem under consideration.
@@
The validity of control (26) relies on the condition that $|\mathcal{E}(0)|<\delta_{T}$ must be satisfied in the transformed plane (see Theorem 4). It is necessary to map this requirement on the initial position and heading angle of the robot in the actual plane. In this direction, it follows from (50) that the condition
@@ VII-A Numerical Simulations
For illustration, consider the scenario described in Example 1, where we have nonconcentric circles with $\mu=\sqrt{5/2}$ and $\lambda=1/2$. Let the robot in the actual plane be characterized by its initial position and heading angle $[x(0),y(0)]^{\top}=[-0.9,-0.6653]^{\top}$ and $\theta(0)=-60^{\circ}$, respectively, and move with constant linear speed $v=1$ m/s. It can be obtained that $|r(0)|=1.1192$ and $\phi(0)=-143.5274^{\circ}$. For $\alpha=\alpha_{s}=0.5$, $\delta_{T}=0.1325$ using (21). It can be verified that (66) is satisfied, since $\eta_{a}(0)=-0.045$, $\eta_{b}(0)=-0.0511$, and $\eta(0)=0.6428$, with $\operatorname{sgn}(\Delta)=1$ and $\sigma=0.5$, resulting in $(\eta_{a}^{2}(0)+\eta_{b}^{2}(0)=0.0046)<(\delta_{T}^{2}\eta^{2}(0)=0.0072)$. Using (17) and (39), the initial position and heading angle in the transformed plane are obtained as $[\rho_{x}(0),\rho_{y}(0)]^{\top}=[0.0016,-0.6039]^{\top}$ and $\gamma(0)=2.3326^{\circ}$, respectively. It can be further verified that $(|\mathcal{E}(0)|=0.106)<(\delta_{T}=0.1325)$.
@@
We conducted simulations using control law (6) with gain $\kappa=0.02$ and plotted various attributes in the actual plane as shown in Fig. 6. It is evident from Fig. 6(a) that the robot’s trajectory converges to the desired circular orbit $\mathcal{C}$ while being confined within the nonconcentric outer boundary $\mathcal{C}^{\prime}$. The absolute value of error (16) converges to the origin, as depicted in Fig. 6(b). The control $\omega$ with roots $\alpha_{s}$ and $\alpha_{\ell}$ is plotted in Fig. 6(c) and Fig. 6(d), respectively. For both values of $\alpha$, $\omega$ converges to $1$, as desired. Further, Fig. 7 illustrates the robot’s motion in the transformed plane and the results are as expected. Please note the difference in the trajectory plots Fig. 7(a) (trajectory-constraining) and Fig. 7(d) (obstacle-avoidance) obtained for $\alpha_{s}$ and $\alpha_{\ell}$, respectively. The robot’s motion is in the anticlockwise direction in Fig. 7(a), while it is in the clockwise direction in Fig. 7(d), in accordance with Theorem 3.
@@ VII-B Experimental Results
To evaluate the performance of the proposed controller (6) in real-world scenarios, we conducted experiments using a Khepera IV111k-team.com/khepera-iv differential drive ground robot. The experiment was facilitated by a motion capture (MoCap) validation system comprising overhead cameras. The MoCap system used Qualisys Track Manager (QTM) software for data recording and processing, providing precise feedback on the robot’s pose data, including position and heading angle. Control commands were sent to the Khepera IV robot based on measurements obtained from QTM using the Robot Operating System (ROS).
To accommodate the size of Khepera IV robot, we adjusted the initial conditions slightly. The robot commenced its motion from the position $[x(0),y(0)]^{\top}=[-0.27\text{\leavevmode\nobreak\ m},-1.15\text{\leavevmode\nobreak\ m}]^{\top}$ with an initial heading angle of $\theta(0)=-30^{\circ}$. Given that the Khepera IV is a differential drive robot, we convert the linear and angular velocities of the unicycle model to the individual wheel velocities of the Khepera IV robot as follows:
$v_{r}=v+\left(\frac{d_{w}}{2}\right)\omega,\qquad v_{\ell}=v-\left(\frac{d_{w}}{2}\right)\omega,$ (81)
where $v_{r}$ and $v_{\ell}$ are the speeds of the right and left wheels, respectively, and $d_{w}=10.54$ cm is the distance between the two wheels of the Khepera IV robot. The robot was operated at a constant linear speed of $v=0.5$ m/s, adhering to its hardware limit for each wheel speed of $0.814$ m/s. Using the control law (6), we derived $v_{r}$ and $v_{\ell}$ in (81). Fig. 8(a) shows the top view of the experimental result at $t=21$ seconds where the red dotted line depicts the robot’s path. Throughout the experiment, the robot’s trajectory was recorded using the QTM software, as depicted in Fig.8(b). Here, the $X$-axis and $Y$-axis denote the real and imaginary axes, respectively. Further, Fig. 8(c) plots the control law applied during the experiment, demonstrating that the turn rate converges to $\omega\approx 0.5$ rad/sec at the steady state.
@@
VIII Conclusion and Further Remarks
Appendix A Proof of Theorem 3
@@
Here, Case (I) and Case (III) refer to $\alpha=\alpha_{s}$ and the rest two cases are for $\alpha=\alpha_{\ell}$. This concludes the proof. ∎
Appendix B Proof of Lemma 5
@@
which, upon multiplication by ${\rm e}^{-i\phi}$ on both sides, yields
@@
where $\vartheta_{1}$ and $\vartheta_{2}$ can be simplified using (41) as: